Yasmin Corporation is comparing two different capital structures, an all-equity plan (Plan I) and a levered plan (Plan II). Under Plan I, Yasmin would have 185,000 shares of stock outstanding. Under Plan II, there would be 135,000 shares of stock outstanding and $2.7 million in debt outstanding. The interest rate on the debt is 5 percent and there are no taxes.

c.

What is the break-even EBIT? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars (e.g., 1,234,567).)

  

  Break-even EBIT $   
Answer #1
At breakeven point the Earnings Per Share (EPS) of both the plans will be equal
EPS of Plan I = EPS of Plan II
EBIT / Number of shares = (EBIT - Interest) / Number of shares
EBIT / 185000 = (EBIT - 2700000*5%) / 135000
EBIT / 185000 = (EBIT - 135000) / 135000
135000 * EBIT / 185000 = EBIT - 135000
135/185*EBIT = EBIT - 135000
135000 = EBIT - 135/185*EBIT
135000 = (185-135)/185*EBIT
135000 = 50/185*EBIT
EBIT = 135000*185/50 499500
break-even EBIT 499500